A square's side is increased by 2 yards, which corresponds to an increase in the area by 56 square yards. How many yards is the side of the initial square?

Answer by mananth(16946)

original side length = x
new length = (x+2)
(x+2)^2= x^2+56
x^2+4x+4 =x^2+56
4x+4 =56
4x=52
x=13 yards ---original length of side